# Making Offline RL Online: Collaborative World Models for Offline Visual Reinforcement Learning

### v3

@@ 4 Experiments
In this section, we present (i) quantitative comparisons with existing visual RL algorithms; (ii) discussions on the influence of discrepancies between the source and target domains; (iii) ablation studies of each proposed training stage; and (iv) further analyses of value overestimation.
@@ 4.1 Experimental Setups
We evaluate CoWorld across three visual control environments, i.e., Meta-World (Yu et al., 2019), RoboDesk (Kannan et al., 2021), and DeepMind Control Suite (DMC) (Tassa et al., 2018), including both cross-task and cross-environment setups (Meta-World $\rightarrow$ RoboDesk). Inspired by D4RL (Fu et al., 2020), we build offline datasets of medium-replay quality using DreamerV2 (Hafner et al., 2021). The datasets comprise all the samples in the replay buffer collected during the training process until the policy attains medium-level performance, defined as achieving $1/3$ of the maximum score that the DreamerV2 agent can achieve. Please refer to Appendix B.2 for further results of CoWorld trained with medium-expert offline data.
@@
We compare CoWorld with both model-based and model-free RL approaches, including Offline DV2 (Lu et al., 2023), DrQ+BC (Lu et al., 2023), CQL (Lu et al., 2023), CURL (Laskin et al., 2020), and LOMPO (Rafailov et al., 2021). In addition, we introduce the DV2 Finetune method, which involves taking a DreamerV2 (Hafner et al., 2021) model pretrained in the online source domain and subsequently finetuning it in the offline target dataset. Furthermore, DV2 Finetune can be integrated with the continual learning method, Elastic Weight Consolidation (EWC) (Kirkpatrick et al., 2017), to regularize the model for preserving source domain knowledge, i.e., Finetune+EWC. Please refer to Appendix D for more details.
@@ 4.2 Cross-Task Experiments on Meta-World
Setup.
Meta-World is an open-source simulated benchmark designed for solving a wide range of robot manipulation tasks. We select $6$ tasks to serve as either the offline dataset or potential candidates for the online auxiliary domain. These tasks include: Door Close ($\textbf{DC}^{*}$), Button Press (BP), Window Close (WC), Handle Press (HP), Drawer Close (DC), Button Topdown (BT).
Comparisons with offline visual RL methods.
As shown in Table 2, we compare the results of CoWorld with other models on Meta-World. CoWorld achieves the best performance in all $6$ tasks. Notably, it outperforms Offline DV2 (Lu et al., 2023), a method also built upon DreamerV2 and specifically designed for offline visual RL.
Comparisons with online-to-offline finetuning.
In Table 2, DV2 Finetune achieves the second-best results by leveraging transferred knowledge from the auxiliary source domain. However, we observe that its performance experiences a notable decline in scenarios (e.g., Meta-World $\rightarrow$ RoboDesk) involving significant data distribution shifts between the source and the target domains in visual observation, physical dynamics, reward definition, or even the action space of the robots. Another important baseline model is DV2 Finetune+EWC, which focuses on mitigating the catastrophic forgetting of the knowledge obtained in source domain pretraining. Nevertheless, without additional model designs for domain adaptation, retaining source domain knowledge may eventually lead to a decrease in performance in the target domain. Moreover, it is interesting to observe that the LOMPO model suffers from the negative transfer effect when incorporating a source pretraining stage. It achieves an average return of $1{,}712$ when it is trained from scratch in the offline domain while achieving an average return of $792$ for online-to-offline finetuning. It implies that a naïve transfer learning method may degenerate the target performance by introducing unexpected bias.
Figure 3: Left: The value in each grid signifies the ratios of returns achieved by CoWorld compared to Offline DV2. Highlighted grids represent the top-performing source domain. Right: Target returns on Drawer Close (DC*) with different source domains. Multi-source CoWorld adaptively selects a useful source (Door Close) and achieves comparable results with the top-performing single-source CoWorld.
@@
(a) Meta-World $\rightarrow$ RoboDesk (Push Green Button)
(b) Y-axis $=$ Estimated value -- True value
Results with a random source domain.
One may cast doubt on the influence of domain discrepancies between the auxiliary environment and the target offline dataset. In Figure 3 (Left), the transfer matrix of CoWorld among the $6$ tasks of Meta-World is presented, where values greater than $1$ indicate positive domain transfer effects. Notably, there are challenging cases with weakly related source and target tasks. In the majority of cases ($26$ out of $30$), CoWorld outperforms Offline DV2, as illustrated in the heatmap.
Results with multiple source domains.
4.3 Cross-Environments: Meta-World to RoboDesk
Setup.
To explore cross-environment domain transfer, we employ four tasks from RoboDesk to construct individual offline datasets, as specified in Figure 4. These tasks require handling randomly positioned objects with image inputs. Table 1 presents the distinctions between the two environments in physical dynamics, action space, reward definitions, and visual appearances. For the best-source experiments, we manually select one source domain from Meta-World. For the multi-source experiments, we jointly use all aforementioned Meta-World tasks as the source domains.
Results.
Figure 4 presents quantitative results of CoWorld, where it outperforms Offline DV2 and DV2 Finetune by large margins. In contrast to prior findings, directly fine-tuning the source world model in this cross-environment setup, where there are more pronounced domain discrepancies, does not result in significant improvements in the final performance. In comparison, CoWorld more successfully addresses these challenges by leveraging domain-specific world models and RL agents, and explicitly aligning the state and reward spaces across domains. We also showcase the performance of multi-source CoWorld, which achieves comparable results to the best-source model that exclusively uses our designated source domain.
4.4 Cross-Dynamics Experiments on DMC
Setup.
Results.
@@ 5 Related Work
Learning control policies from images is critical in real-world applications. Existing approaches can be grouped by the use of model-free (Laskin et al., 2020; Schwarzer et al., 2021; Stooke et al., 2021; Xiao et al., 2022; Parisi et al., 2022) or model-based (Hafner et al., 2019, 2020, 2021; Seo et al., 2022; Pan et al., 2022; Hafner et al., 2022; Mazzaglia et al., 2023; Micheli et al., 2023; Zhang et al., 2023; Ying et al., 2023) RL algorithms. In offline RL, agents leverage pre-collected offline data to optimize policies and encounter challenges associated with value overestimation (Levine et al., 2020). Previous methods mainly suggest taking actions that were previously present in the offline dataset or learning conservative value estimations (Fujimoto et al., 2019; Kumar et al., 2020; Chen et al., 2022; Yu et al., 2020, 2021; Rigter et al., 2022). Recent approaches have introduced specific techniques to address the challenges associated with offline visual RL (Mandlekar et al., 2019; Dasari et al., 2019; Levine et al., 2020; Agarwal et al., 2020; Rafailov et al., 2021; Yu et al., 2022; Seo et al., 2022; Zang et al., 2023; Cho et al., 2022; Lu et al., 2023). Rafailov et al. (2021) proposed to handle high-dimensional observations with latent dynamics models and uncertainty quantification. Cho et al. (2022) proposed synthesizing the raw observation data to append the training buffer, aiming to mitigate the issue of overfitting. In a related study, Lu et al. (2023) established a competitive offline visual RL model based on DreamerV2 (Hafner et al., 2021), so that we use it as a significant baseline of our approach. In contrast to previous methods, we innovatively frame offline visual RL as an offline-online-offline transfer learning problem to allow for the use of auxiliary simulators to further mitigate the value overestimation issues.
Transfer RL.
@@
Our work is also related to transfer RL, which is known as to utilize the knowledge learned in past tasks to facilitate learning in unseen tasks (Zhu et al., 2020; Sekar et al., 2020; Zhang et al., 2020; Sun et al., 2021; Zhang et al., 2021; Eysenbach et al., 2021; Yang & Nachum, 2021; Sun et al., 2022; Ghosh et al., 2023; Kumar et al., 2023; Rafailov et al., 2023; Liu et al., 2023; Nakamoto et al., 2023). In the context intersected with visual RL, CtrlFormer (Mu et al., 2022) learns a transferable state representation via a sample-efficient vision Transformer. APV (Seo et al., 2022) executes action-free world model pretraining on source-domain videos and finetunes the model on downstream tasks. Choreographer (Mazzaglia et al., 2023) builds a model-based agent that exploits its world model to learn and adapt skills in imaginations, the learned skills are adapted to new domains using a meta-controller. VIP (Ma et al., 2023) presents a self-supervised, goal-conditioned value-function objective, which enables the use of unlabeled video data for model pretraining. Different from the aforementioned approaches, our study concentrates on performance within the offline domain, in the sense that we provide a pilot exploration of transfer learning for offline visual control.
@@ 6 Conclusions and Limitations
In this paper, we proposed a transfer RL method named CoWorld, which mainly tackles the difficulty in representation learning and value estimation in offline visual RL. The key idea is to exploit accessible online environments to train an auxiliary RL agent to offer additional value assessment. To address the domain discrepancies and to improve the offline policy, we present specific technical contributions of cross-domain state alignment, reward alignment, and min-max value constraint. CoWorld demonstrates competitive results across three RL benchmarks.
An unsolved problem of CoWorld is the increased computational complexity associated with the training phase in auxiliary domains (see Appendix B.5). It is valuable to improve the training efficiency in future research.
7 Broader Impacts
Acknowledgements
This work was supported by NSFC (62250062, U19B2035, 62106144), Shanghai Municipal Science and Technology Major Project (2021SHZDZX0102), the Fundamental Research Funds for the Central Universities, and Shanghai Sailing Program (21Z510202133) from the Science and Technology Commission of Shanghai Municipality.
@@ A.1 Framework of CoWorld
A.2 World Model Learning
We adopt the framework of the world model used in  (Hafner et al., 2021). The image encoder is a Convolutional Neural Network (CNN). The image predictor is a transposed CNN and the transition, reward, and discount factor predictors are Multi-Layer Perceptrons (MLPs). The discount factor predictor serves as an estimate of the probability that an episode will conclude while learning behavior based on model predictions. The encoder and decoder take 64 × 64 images as inputs.
The loss function of the target world model (i.e., Eq. (3)) is jointly minimized with respect to the $\phi^{\prime}$ that contains all parameters of the target world model. The image predictor, reward predictor, discount predictor, and transition predictor are trained to maximize the log-likelihood of their individual targets through the distributions they produce.
@@ A.3 Behavior Learning
For behavior learning of CoWorld, we use the actor-critic learning architecture of DreamerV2 (Hafner et al., 2021). The $\lambda$-target $V_{t}^{(T)}$ in Eq. (6) is defined as follows:
@@
where $\lambda$ is set to 0.95 for considering more on long horizon targets. The actor and critic are both MLPs with ELU (Clevert et al., 2015) activations while the world model is fixed. The target actor and critic are trained with guidance from the source critic, and regress the $\lambda$-return with a squared loss.
The source actor and critic are:
@@
We train the source actor $\pi_{\psi^{\prime}}$ by maximizing an objective function:
$\mathcal{L}(\psi^{\prime})$ $=\ \mathbb{E}_{p_{\phi^{\prime}},p_{\psi^{\prime}}}\Big{[}\sum_{t=1}^{H-1}(\underbrace{\beta\mathrm{H}\left[a_{t}^{(S)}\mid\hat{z}_{t}^{(S)}\right]}_{\text{entropy regularization}}+\underbrace{\rho V_{t}^{(S)}}_{\text{dynamics backprop}}+\underbrace{(1-\rho)\ln\pi_{\psi^{\prime}}(\hat{a}_{t}^{(S)}\mid\hat{z}_{t}^{(S)})\texttt{sg}(V_{t}^{(S)}-v_{\xi^{\prime}}(\hat{z}_{t}^{(S)}))}_{\text{REINFORCE}}\Big{]}.$ (10)
The source critic $v_{\xi^{\prime}}$ is optimized by minimizing a loss function:

### v4

@@ 4.1 Experimental Setups
We evaluate CoWorld across three visual control environments, i.e., Meta-World [54], RoboDesk [18], and DeepMind Control Suite (DMC) [47], including both cross-task and cross-environment setups (Meta-World $\rightarrow$ RoboDesk). Inspired by D4RL [9], we build offline datasets of medium-replay quality using DreamerV2 [16]. The datasets comprise all the samples in the replay buffer collected during the training process until the policy attains medium-level performance, defined as achieving $1/3$ of the maximum score that the DreamerV2 agent can achieve. Please refer to Appendix B.2 for further results of CoWorld trained with medium-expert offline data.
@@
We compare CoWorld with both model-based and model-free RL approaches, including Offline DV2 [25], DrQ+BC [25], CQL [25], CURL [22], and LOMPO [39]. In addition, we introduce the DV2 Finetune method, which involves taking a DreamerV2 [16] model pretrained in the online source domain and subsequently finetuning it in the offline target dataset. Furthermore, DV2 Finetune can be integrated with the continual learning method, Elastic Weight Consolidation (EWC) [19], to regularize the model for preserving source domain knowledge, i.e., Finetune+EWC. Please refer to Appendix E for more details.
@@ 4.2 Cross-Task Experiments on Meta-World
Meta-World is an open-source simulated benchmark designed for solving a wide range of robot manipulation tasks. We select $6$ tasks as either the offline dataset or potential candidates for the online auxiliary domain. These tasks include: Door Close ($\textbf{DC}^{*}$), Button Press (BP), Window Close (WC), Handle Press (HP), Drawer Close (DC), Button Topdown (BT).
Main results.
As shown in Table 2, we compare the results of CoWorld with other models on Meta-World. CoWorld achieves the best performance in all $6$ tasks. Notably, it outperforms Offline DV2 [25], a method also built upon DreamerV2 and specifically designed for offline visual RL. For the online-to-offline finetuning models, DV2 Finetune achieves the second-best results by leveraging transferred knowledge from the auxiliary source domain. However, we observe that its performance experiences a notable decline in scenarios (e.g., Meta-World $\rightarrow$ RoboDesk) involving significant data distribution shifts between the source and the target domains in visual observation, physical dynamics, reward definition, or even the action space of the robots. Another important baseline model is DV2 Finetune+EWC, which focuses on mitigating the catastrophic forgetting of the knowledge obtained in source domain pretraining. Nevertheless, without additional model designs for domain adaptation, retaining source domain knowledge may eventually lead to a decrease in performance in the target domain. The LOMPO model suffers from the negative transfer effect when incorporating a source pretraining stage. It achieves an average return of $1{,}712$ when it is trained from scratch in the offline domain while achieving an average return of $792$ for online-to-offline finetuning. It implies that a naïve transfer learning method may degenerate the target performance due to unexpected bias.
Figure 3: Left: The value in each grid indicates the ratio of returns achieved by CoWorld compared to Offline DV2. Highlighted grids represent the top-performing source domain. Right: Returns on Drawer Close (DC*) with different source domains, where the multi-source CoWorld (yellow line) is shown to automatically discover (i.e., Door Close) as the source domain and achieve comparable results with the top-performing single-source CoWorld (red line).
@@
Results with a random source domain.
Given that we present the best-source results in Table 2, where we manually select one source task from Meta-World, one may cast doubt on the influence of domain discrepancies between the auxiliary environment and the target offline dataset. In Figure 3 (Left), the transfer matrix of CoWorld among the $6$ tasks of Meta-World is presented, where values greater than $1$ indicate positive domain transfer effects. Notably, there are challenging cases with weakly related source and target tasks. In the majority of cases ($26$ out of $30$), CoWorld outperforms Offline DV2, as illustrated in the heatmap.
Results with multiple source domains.
4.3 Cross-Environments: Meta-World to RoboDesk
To explore cross-environment transfer with more significant domain gaps, we employ four tasks from RoboDesk to construct individual offline datasets, i.e., Push Button, Open Slide, Drawer Open, Upright Block off Table. These tasks require handling randomly positioned objects with image inputs. Table 1 presents the differences between the two environments in physical dynamics, action space, reward definitions, and visual appearances.
Figure 4 presents quantitative comparisons, where CoWorld outperforms Offline DV2 and DV2 Finetune by large margins. For the best-source experiments, we manually select one source domain from Meta-World. For the multi-source experiments, we jointly use all Meta-World tasks as the source domains. In contrast to prior findings, directly finetuning the source world model in this cross-environment setup, where there are more pronounced domain discrepancies, does not result in significant improvements in the final performance. In comparison, CoWorld more successfully addresses these challenges by leveraging domain-specific world models and RL agents, and explicitly aligning the state and reward spaces across domains. We also showcase the performance of multi-source CoWorld, which achieves comparable results to the best-source model that exclusively uses our designated source domain.
4.4 Cross-Dynamics Experiments on DMC
(a) Meta-World $\rightarrow$ RoboDesk (Push Green Button)
(b) Y-axis $=$ Estimated value -- True value
@@ 4.5 Further Analyses
Dependence of CoWorld to domain similarities.
We further investigate the dependence of CoWorld on domain similarity from the perspectives of different observation spaces and reward spaces. We first explore how CoWorld performs when we only have source domains with significantly distinct observation spaces from the target domain. As illustrated in Table 4, the agent receives low-dimensional state inputs in the source domain (Meta-World) and high-dimensional images in the target domain (RoboDesk). We can see that CoWorld outperforms Offline DV2 by $13.3\%$ and $34.0\%$ due to the ability to leverage low-dimensional source data effectively. Notably, the finetuning method (DV2 Finetune) is not applicable in this scenario. In Table 5, we also observe that CoWorld benefits from a source domain, even with a significantly different reward signal. Unlike previous experiments, we use a sparse reward function for the source Meta-World tasks. It is set to $500$ only upon task completion and remains 00 before that. The experimental results demonstrate that although excessively sparse rewards can hinder the training process, CoWorld still achieves an average performance gain of $6.6\%$ compared to DV2 Finetune under the same setting.
Comparison to jointly training one world model across domains.
Notably, CoWorld is implemented with separate world models for the source and target domains. Alternatively, we can employ a jointly trained world model across various domains for more efficient memory usage. In Table 6, we compare the results from the original CoWorld and “Multi-Task DV2”. Multi-Task DV2 involves training DreamerV2 on both offline and online data with a joint world model and separate actor-critic models. CoWorld consistently performs better. Intuitively, using separate world models allows the source and target domains to have different physical dynamics, observation spaces, or reward formations, as the scenarios shown in Table 4 and Table 5.
Table 4: Experiments with significantly distinct observation spaces across domains. We use low-dimensional state data as inputs for the RL agents in the source domain and high-dimensional image observations in the target domain. MW represents Meta-World and RD stands for RoboDesk.
Method MW: Button Press $\rightarrow$ RD: Push Button MW: Window Close $\rightarrow$ RD: Open Slide
Offline DV2 347 $\pm$ 24 156 $\pm$ 46
CoWorld 393 $\pm$ 64 209 $\pm$ 43
Table 5: Experiments with significantly distinct reward formations across domains. We use sparse rewards in the source domain while maintaining the dense rewards in the target domain.
Method MW: Button Press $\rightarrow$ RD: Push Button MW: Window Close $\rightarrow$ RD: Open Slide
DV2 Finetune 314 $\pm$ 51 173 $\pm$ 39
CoWorld 335 $\pm$ 28 184 $\pm$ 32
Table 6: Comparison to jointly training one world model across domains (Multi-Task DV2).
Method MW: Button Press $\rightarrow$ RD: Push Button MW: Window Close $\rightarrow$ RD: Open Slide
Multi-Task DV2 342 $\pm$ 29 173 $\pm$ 22
CoWorld 428 $\pm$ 42 202 $\pm$ 19
@@ 5 Related Work
Learning control policies from images is critical in real-world applications. Existing approaches can be grouped by the use of model-free [22, 41, 44, 48, 36] or model-based [15, 14, 16, 43, 35, 13, 28, 29, 61, 51] RL algorithms. In offline RL, agents leverage pre-collected offline data to optimize policies and encounter challenges associated with value overestimation [23]. Previous methods mainly suggest taking actions that were previously present in the offline dataset or learning conservative value estimations [11, 21, 4, 55, 53, 40]. Recent approaches have introduced specific techniques to address the challenges associated with offline visual RL [27, 7, 23, 2, 39, 52, 43, 57, 5, 25]. Rafailov et al. [39] proposed to handle high-dimensional observations with latent dynamics models and uncertainty quantification. Cho et al. [5] proposed synthesizing the raw observation data to append the training buffer, aiming to mitigate the issue of overfitting. In a related study, Lu et al. [25] established a competitive offline visual RL model based on DreamerV2 [16], so that we use it as a significant baseline of our approach.
@@
Our work is also related to transfer RL, which is known as to utilize the knowledge learned in past tasks to facilitate learning in unseen tasks [64, 42, 58, 45, 59, 8, 49, 46, 12, 20, 38, 24, 33]. Most existing approaches related to offline dataset + simulator focus on the offline-to-online setup, where the policy is initially pretrained on the offline dataset and then finetuned and deployed on an interactive environment [33, 60, 56, 63]. These methods aim to bridge the gap between offline and online learning and facilitate fast adaptation of the model to the online environment. In contrast, we explore the online-to-offline setup, which provides a new remedy for the value over-estimation problem. Additionally, Niu et al. [34] introduces a dynamics-aware hybrid offline-and-online framework to integrate offline datasets and online simulators for policy optimization. Unlike CoWorld, this method primarily focuses on low-dimensional MDPs and cannot be directly used in visual control tasks. In the context of visual RL, CtrlFormer [31] learns a transferable state representation via a sample-efficient vision Transformer. APV [43] executes action-free world model pretraining on source-domain videos and finetunes the model on downstream tasks. Choreographer [28] builds a model-based agent that exploits its world model to learn and adapt skills in imaginations, the learned skills are adapted to new domains using a meta-controller. VIP [26] presents a self-supervised, goal-conditioned value-function objective, which enables the use of unlabeled video data for model pertaining. Unlike previous methods, we handle offline visual RL using auxiliary simulators, mitigating the value overestimation issues with co-trained world models.
@@ 6 Conclusions and Limitations
In this paper, we proposed a transfer RL method named CoWorld, which mainly tackles the difficulty in representation learning and value estimation in offline visual RL. The key idea is to exploit accessible online environments to train an auxiliary RL agent to offer additional value assessment. To address the domain discrepancies and to improve the offline policy, we present specific technical contributions of cross-domain state alignment, reward alignment, and min-max value constraint. CoWorld demonstrates competitive results across three RL benchmarks. An unsolved problem of CoWorld is the increased computational complexity associated with the training phase in auxiliary domains (see Appendix B.7). It is valuable to improve the training efficiency in future research.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (Grant No. 62250062, 62106144, 62302246), the Shanghai Municipal Science and Technology Major Project (No. 2021SHZDZX0102), the Fundamental Research Funds for the Central Universities, and the CCF-Tencent Rhino-Bird Open Research Fund. This work was also supported by the Natural Science Foundation of Zhejiang Province, China (No. LQ23F010008), the High-Performance Computing Center at Eastern Institute of Technology (Ningbo), and the Ningbo Institute of Digital Twin.
Appendix
In this appendix, we provide the following supplementary materials: (A) Details of the proposed model, including further descriptions of the learning schemes, the notations, the world model architecture, the behavior learning objective functions, and hyperparameters. (B) Additional experimental results, including visualization of the learned policy, quantitative results on offline datasets with mixed data quality, comparison to using a pre-trained foundation model such as R3M, and computational efficiency. (C) Implementation details of the multi-source CoWorld model and further empirical analysis on the selected source domain. (D) Detailed setups of the source and target domains. (E) Details of the compared methods. (F) Potential social impacts of the proposed method.
@@ A.1 Framework of CoWorld
For notations, we use the superscript $S$ and $T$ to represent data from the source and target domains. Additionally, subscripts $(\phi^{\prime},\psi^{\prime},\xi^{\prime})$ and $(\phi,\psi,\xi)$ are employed to distinguish model parameters for different domains. The notations of source and target domains are summarised in Table 7.
Table 7: Notations of the source and target domains.
Domains Model Parameters Data
Source/Online ($S$) World model $\phi^{\prime}$, Actor $\psi^{\prime}$, Critic $\xi^{\prime}$ Raw data $(o_{t}^{(S)},a_{t}^{(S)},r_{t}^{(S)})$, Relabeled reward with infomation from both domains $\tilde{r}_{t}^{(S)}$
Target/Offline ($T$) World model $\phi$, Actor $\psi$, Critic $\xi$ Raw data $(o_{t}^{(T)},a_{t}^{(T)},r_{t}^{(T)})$
A.2 World Model
We adopt the framework of the world model used in  [16]. The image encoder is a Convolutional Neural Network (CNN). The image predictor is a transposed CNN and the transition, reward, and discount factor predictors are Multi-Layer Perceptrons (MLPs). The discount factor predictor serves as an estimate of the probability that an episode will conclude while learning behavior based on model predictions. The encoder and decoder take $64\times 64$ images as inputs.
@@ A.3 Behavior Learning
For the behavior learning of CoWorld, we use the actor-critic method from DreamerV2 [16]. The $\lambda$-target $V_{t}^{(T)}$ in Eq. (6) is defined as follows:
@@
where $\lambda$ is set to $0.95$ for considering more on long horizon targets. The actor and critic are both MLPs with ELU activations [6]. The target actor and critic are trained with guidance from the source critic and regress the $\lambda$-return with a squared loss. The world model is fixed during behavior learning. The source actor and critic are:
@@
We train the source actor $\pi_{\psi^{\prime}}$ by maximizing
$\mathcal{L}(\psi^{\prime})=\ \mathbb{E}_{p_{\phi^{\prime}},p_{\psi^{\prime}}}\Big{[}\sum_{t=1}^{H-1}(\underbrace{\beta\mathrm{H}\left[a_{t}^{(S)}\mid\hat{z}_{t}^{(S)}\right]}_{\text{entropy regularization}}+\underbrace{\rho V_{t}^{(S)}}_{\text{dynamics backprop}}$ (10)
$+\underbrace{(1-\rho)\ln\pi_{\psi^{\prime}}(\hat{a}_{t}^{(S)}\mid\hat{z}_{t}^{(S)})\texttt{sg}(V_{t}^{(S)}-v_{\xi^{\prime}}(\hat{z}_{t}^{(S)}))}_{\text{REINFORCE}}\Big{]}.$
The source critic $v_{\xi^{\prime}}$ is optimized by minimizing